OpenStudy (anonymous):

solve the equation then check your solution a+8=4 ________ -2.1

mathslover (mathslover):

is this the question : \[\huge{\frac{a+8}{-2.1}=4}\]

OpenStudy (anonymous):

i got 8.4 when i tryed

mathslover (mathslover):

sorry u r not exactly right : \[\huge{a+8=-8.4}\] can u solve further ?

mathslover (mathslover):

Just remember : \[\huge{\frac{a}{b}=c}\] \[\huge{a=bc}\]

OpenStudy (anonymous):

-4

mathslover (mathslover):

may i know that what is ? \[\huge{-8.4-8=?}\]

OpenStudy (anonymous):

-16.4

mathslover (mathslover):

so here you go with the correct answer

mathslover (mathslover):

remember : \[\huge{a+b=c}\]then \[\huge{a=c-b}\]

OpenStudy (anonymous):

thats not one of my answers i can choose from a. -4 b. -25.2 c. -25.1 d. 8.4

OpenStudy (anonymous):

its d right

mathslover (mathslover):

is this the question ? : \[\huge{\frac{a+8}{-2.1}=4}\] ?????

OpenStudy (ganpat):

i guess, the bottom part -2.1... would be some different question :-p..